# CoAM: Corpus of All-Type Multiword Expressions

### v2

@@ 1 Introduction
The results reveal that a fine-tuned Qwen model with 72B parameters greatly outperforms MaW, demonstrating the effectiveness of LLM fine-tuning. On the other hand, all approaches suffer from low recall (e.g., 50.7% for Qwen-72B). Further analysis shows that fine-tuned LLMs struggle more with detecting Noun and Clause MWEs than with detecting Verb MWEs. MWEs not contained in WordNet Miller (1995), e.g., real estate, were found particularly difficult to identify, presumably because they are less widely recognized as MWEs.
@@ 4.1 Data Selection and Preprocessing
For all sources, we took the first 10 (or all, when there are less than 10) sentences from each article or talk and presented them in original order. Each sentence was tokenized using spaCy (we use version 3.7.1 and the en_core_web_trf model throughout this paper), with the exception of the UD sentences, which were already tokenized. Note that what we call words are tokens given in this manner.
@@ 4.5 Inter-Annotator Agreement
With two annotators being assigned to each sentence, we measured the inter-annotator agreement (IAA) by the MWE-based (exact-match) F1 score. The IAA was found to be only 37.3%, which could be attributed to annotators failing to annotate valid MWEs. We thus suggest that MWE datasets should be constructed with at least two independent annotators to reduce missed MWEs.
@@ 5 MWEI Approaches
We use CoAM to evaluate the following MWEI approaches.
@@ 6.2 Results and Analysis
Given the low recall across all systems, we conduct further analysis to determine which MWEs the models particularly struggle to identify.
@@
Ablation Study on LLM Fine-Tuning
@@
As shown in Table 6a, fine-tuned Qwen-72B outperforms the best MaW system in all of the F1 score, precision, and recall. However, the LLM takes substantial memory and compute, and MaW is more efficient in this regard. MaW’s encoder model (bert-base-uncased) has only 110M parameters, taking approximately $2\times 110\textrm{M}\times 4=880\textrm{M}$ bytes in total ($2$ is the number of encoders). 4-bit quantized Qwen-72B takes $72,000\textrm{M}\times 0.5=36,000\textrm{M}$ bytes, taking roughly 40 times more space than MaW. On the other hand, the performance of 4-bit quantized Qwen-7B, which takes roughly 4 times more space, is comparable with Rule+DCA.
@@ 7 Conclusion
In this paper, we constructed CoAM, a high-quality dataset of 1.3K sentences for MWEI covering all types of MWEs. Using a combination of human review and automated consistency checking, we addressed consistency issues that have been a problem for previous MWE datasets, enabling more accurate evaluation results for future work in MWEI. We used CoAM to evaluate two MWEI approaches: MaW and LLM fine-tuning. Our largest fine-tuned LLM performed the best, but all systems suffered from low recall. Consequently, we argue that MWEI overall remains a challenging task.
@@ Acknowledgments
We are grateful for the insightful comments by Yuki Arase, Shohei Higashiyama, and anonymous reviewers. To polish the content of this paper, we used ChatGPT111111 https://chatgpt.com as a writing assistant tool. Financially, this work was supported by JST SPRING Grant Number JPMJSP2140.

### v3

@@ 1 Introduction
The results reveal that the fine-tuned Qwen 72B model greatly outperforms MaW, demonstrating the effectiveness of LLM fine-tuning. On the other hand, all approaches suffer from low recall (e.g., 52.8% for Qwen-72B). Further analysis shows that fine-tuned LLMs struggle more with detecting Noun and Clause MWEs than with detecting Verb MWEs. MWEs not contained in WordNet Miller (1995), e.g., real estate, were found particularly difficult to identify, presumably because they are less widely recognized as MWEs.
@@ 4.1 Data Selection and Preprocessing
For all sources, we took the first 10 (or all, when there are less than 10) sentences from each article or talk and presented them in original order. Each sentence was tokenized using spaCy (we use version 3.7.1 and the en_core_web_trf model throughout this paper), with the exception of the UD sentences, which were already tokenized. Note that we call tokens given in this manner words.
@@ 4.5 Inter-Annotator Agreement
With two annotators being assigned to each sentence, we measured the inter-annotator agreement (IAA) by the MWE-based (exact-match) F1 score. The IAA was found to be only 37.3%, which could be attributed to annotators failing to annotate valid MWEs. We thus suggest that MWE datasets be constructed with at least two independent annotators to reduce missed MWEs and that the annotators be adequately trained to ensure they fully understand the annotation guidelines.
@@ 5 MWEI Approaches
We use CoAM to evaluate two MWEI approaches.
@@ 6.2 Results and Analysis
Given the low recall across all systems, we conduct further analysis to examine which MWEs are particularly difficult to identify.
@@
Recall by Continuity Recall by In WN or Not
F1 P R Continuous Discontinuous      True False
Human 56.5 64.5 51.0 52.1 41.8      56.8 47.0
Table 8: Mean human agreement with the final CoAM annotations. Both train and test sets are included in the data.
@@
As shown in Table 6a, fine-tuned Qwen-72B outperforms the best MaW system in all of the F1 score, precision, and recall. However, MaW is more efficient than the LLM in terms of memory and compute. MaW’s encoder model (bert-base-uncased) has only 110M parameters, taking approximately $2\times 110\textrm{M}\times 4=880\textrm{M}$ bytes in total ($2$ is the number of encoders). 4-bit quantized Qwen-72B takes $72,000\textrm{M}\times 0.5=36,000\textrm{M}$ bytes, taking roughly 40 times more space than MaW. On the other hand, the performance of 4-bit quantized Qwen-7B, which takes roughly 4 times more space, is comparable with Rule+DCA.
7 Discussion
How do the errors made by the automated approaches contrast with human annotators’ errors? To address this question, we perform an analysis using the final CoAM annotations as gold annotations and the raw human annotations as predictions. This is not a completely fair setting for evaluating the humans’ capability of MWEI, as the human annotations partially contribute to the gold standard; however, annotations by an annotator are only one factor in determining the final gold annotations. The gold annotations on each sentence are also affected by another annotator, a reviewer, and the consistency checking process. We thus expect this analysis to shed light on the characteristics of MWEI performance for humans and the systems.
Table 8 shows the human agreement with the gold annotations. Here we find all of F1, precision, and recall are similar to those of fine-tuned Qwen-72B; on average, human recall is only around 50%. However, closer analysis reveals this is not true for all the annotators. The top annotator achieved an F1 of 72.0% with a recall of 71.9%, while the runner-up annotator reached an F1 of 66.0% with a recall of 59.4%.111111 It should be noted that the annotators worked on different sentences, but we assume the annotation difficulty varied minimally. The top annotator holds a linguistics-related degree, and the second has authored a paper on MWEI, which may have served as relevant training for MWEI annotation. These results suggest that automated systems could also achieve higher MWEI performance if the underlying models were better trained.
In Section 6.2, we discussed that discontinuous MWEs remain a key challenge for the MWEI systems. In contrast, Recall by Continuity in Table 8 shows that discontinuous MWEs are not especially challenging for humans, compared to continuous MWEs. This indicates that better modeling for discontinuous MWEs could greatly improve the systems’ performance on this category of MWEs.
@@ 8 Conclusion
In this paper, we constructed CoAM, a high-quality dataset of 1.3K sentences for MWEI. Using a combination of human review and automated consistency checking, we addressed consistency issues that have been a problem for previous MWEI datasets, enabling more accurate evaluation results. We used CoAM to evaluate two MWEI approaches: MaW and LLM fine-tuning. Our largest fine-tuned LLM performed the best, but all systems suffered from low recall. Consequently, we argue that MWEI overall remains a challenging task.
@@ Acknowledgments
We are grateful for the insightful comments by Yuki Arase, Shohei Higashiyama, and anonymous reviewers. To polish the content of this paper, we used ChatGPT121212 https://chatgpt.com as a writing assistant tool. Financially, this work was supported by JST SPRING Grant Number JPMJSP2140.